# Soft symmetries of topological orders

### v3

@@
[…]
eqs
$\BODY\BODY\begin{split}\BODY\end{split}$ (1)
@@
(2+1)D topological orders possess emergent symmetries given by a group $\text{Aut}(\mathcal{C})$, which consists of the braided tensor autoequivalences of the modular tensor category $\mathcal{C}$ that describes the anyons. In this paper we discuss cases where $\text{Aut}(\mathcal{C})$ has elements that neither permute anyons nor are associated to any symmetry fractionalization but are still non-trivial, which we refer to as soft symmetries. We point out that one can construct topological defects corresponding to such exotic symmetry actions by decorating with a certain class of gauged SPT states that cannot be distinguished by their torus partition function. This gives a physical interpretation to work by Davydov on soft braided tensor autoequivalences. This has a number of important implications for the classification of gapped boundaries, non-invertible spontaneous symmetry breaking, and the general classification of symmetry-enriched topological phases of matter. We also demonstrate analogous phenomena in higher dimensions, such as (3+1)D gauge theory with gauge group given by the quaternion group $Q_{8}$.
@@ II.1 Review: Unitary modular tensor categories
where $\mu=1,\ldots,N_{ab}^{c}$, $d_{a}$ is the quantum dimension of $a$, and the factors $\left(\frac{d_{c}}{d_{a}d_{b}}\right)^{1/4}$ are a normalization convention for the diagrams. The set of anyons contains the identity or vacuum charge 00, and each anyon $a$ has a dual anti-particle $\bar{a}$ such that $a\times\bar{a}=0+\cdots$.
@@
where $\gamma_{a},\gamma_{b},\gamma_{c}$ are phase factors close to unity. Taking equivalence classes of braided tensor autoequivalences modulo natural isomorphisms defines a finite group, denoted Aut$(\mathcal{C})$, which is the group of braided auto-equivalences of $\mathcal{C}$.
@@ II.3 Symmetry action on genus $g$ surfaces
The $U$ symbols characterize the action of $[\varphi]\in\text{Aut}(\mathcal{C})$ on the space of states on a closed genus $g$ surface. We can pick an orthonormal basis of states on the genus $g$ surface as
@@
It is straightforward to verify that in general the symmetry action on every basis state is unchanged by a natural isomorphism. Therefore, the symmetry action only depends on the equivalence class $[\varphi]$. This implies that if a braided autoquivalence $\varphi$ acts as the identity on the space of states on a torus while it acts non-trivially on surfaces of genus $g\geq 2$, then it must be a non-trivial soft symmetry.
@@
The group of braided tensor autoequivalences defines an element of a 3rd cohomology group, $[\mathcal{O}_{3}]\in\mathcal{H}^{3}(\text{Aut}(\mathcal{C}),\mathcal{A})$, where $\mathcal{A}$ is a finite Abelian group defined by fusion of the Abelian anyons in $\mathcal{C}$ [40, 3]. Formulas to compute $[\mathcal{O}_{3}]$ from the $U$-symbols were presented in [3]. Graphically, we can understand $[\mathcal{O}_{3}]$ as a failure of associativity for fusing the codimension-1 sheets. Consider three symmetries $[\varphi_{1}]$, $[\varphi_{2}]$, $[\varphi_{3}]$ $\in\text{Aut}(\mathcal{C})$ and the junctions formed by fusing the sheets as $\varphi_{1}\circ(\varphi_{2}\circ\varphi_{3})$ compared with $(\varphi_{1}\circ\varphi_{2})\circ\varphi_{3}$. As we deform from the first order to the second, an Abelian anyon $O_{3}([\varphi_{1}],[\varphi_{2}],[\varphi_{3}])\in\mathcal{A}$ is emitted from the resulting transition point in space-time. See Fig. 2 for an illustration. We see how $[\mathcal{O}_{3}]$ can be viewed as an obstruction to associativity of the symmetry operations [3].
The data $(\text{Aut}(\mathcal{C}),\mathcal{A},[\mathcal{O}_{3}])$ defines a 2-group $\underline{\text{Aut}(\mathcal{C})}$ that acts on the UMTC $\mathcal{C}$ [3, 41]. An example of a theory with non-trivial $[\mathcal{O}_{3}]$ was discussed in [3, 42], with generalizations where the symmetry is anti-unitary [43].
@@ II.5 Symmetry fractionalization
where $\varphi_{jk}$ denotes $\varphi_{i}\circ\varphi_{j}$, and $M_{xa}$ denotes the $U(1)$ phase resulting from a double braid of $x$ with an Abelian anyon $a\in\mathcal{A}$. If $[\mathcal{O}_{3}]$ is trivial in group cohomology $\mathcal{H}^{3}(\text{Aut}(\mathcal{C}),\mathcal{A})$ twisted by the action of $\text{Aut}(\mathcal{C})$ on $\mathcal{A}$, then there are inequivalent choices of the $\eta$ symbols, which form a torsor over the twisted cohomology $\mathcal{H}^{2}(\text{Aut}(\mathcal{C}),\mathcal{A})$.
The $\eta$-symbols are referred to as the symmetry fractionalization data of the theory, because they specify how the anyons carry fractional quantum numbers under the symmetry group elements. For example, they can be used to characterize the fractional electric charge of anyons in the FQH effect or fractional spin of spinons in quantum spin liquids [44]. Assuming the $[\mathcal{O}_{3}]$ is trivial in $\mathcal{H}^{3}(\text{Aut}(\mathcal{C}),\mathcal{A})$, the $\eta$ symbols further determine an ’t Hooft anomaly $[\mathcal{O}_{4}]\in\mathcal{H}^{4}(\text{Aut}(\mathcal{C}),U(1))$ [40, 3, 45, 46, 47].
@@ II.6 Global symmetry $G$ of quantum many-body system
Consider a quantum many-body system with a global 00-form symmetry group $G$. If the system forms a gapped phase with anyons described by a UMTC $\mathcal{C}$, then the possible symmetry-enriched topological phases will be partially characterized by a group homomorphism
$[\rho]:G\rightarrow\text{Aut}(\mathcal{C}).$ (22)
@@
The group homomorphism $[\rho]$ can be pulled back to determine a cohomology class $[\rho]^{*}[\mathcal{O}_{3}]\in\mathcal{H}^{3}(G,\mathcal{A})$, which is an obstruction for the $\eta$ symbols of $G$ to satisfy associativity. Let us assume that $[\rho]^{*}[\mathcal{O}_{3}]$ is trivial.
@@
where $[{\bf t}]\in\mathcal{H}^{2}(G,\mathcal{A})$. Therefore, the possible choices of symmetry fractionalization for $G$ factor through those for Aut$(\mathcal{C})$, with an additional freedom specified by $[\bf t]$. The above data can be viewed as specifying a 2-homomorphism from $G$, viewed as a 2-group $\underline{G}$ with trivial $1$-form symmetry, to $\underline{\text{Aut}(\mathcal{C})}$:
$\underline{\rho}:\underline{G}\rightarrow\underline{\text{Aut}(\mathcal{C})}.$ (25)
If $[\rho]^{*}[\mathcal{O}_{3}]$ is non-trivial, then the theory is incompatible with an ordinary global $G$ symmetry. Instead we must consider a system with a non-trivial 2-group symmetry $\underline{G}=(G,\mathcal{A},[\rho],[\rho]^{*}[\mathcal{O}_{3}])$. In this sense $[\rho]^{*}[\mathcal{O}_{3}]$ is an obstruction to having the symmetry be an ordinary group $G$.
As mentioned above, the symmetry fractionalization data determines an element $[\mathcal{O}_{4}]\in\mathcal{H}^{4}(\text{Aut}(\mathcal{C}),U(1))$, which in the field theory literature is referred to as the ’t Hooft anomaly. This induces an ’t Hooft anomaly for $G$ via the pullback $[\mathcal{O}_{4,G}]=[\rho]^{*}[\mathcal{O}_{4}]\times\mathcal{O}_{4,\text{rel}}[{\bf t}]$, where $\mathcal{O}_{4,\text{rel}}[{\bf t}]$ is an additional contribution, referred to as a relative anomaly [46], determined by the choice $[\bf t]$, along with the $\eta$ and $U$ symbols for $\text{Aut}(\mathcal{C})$. General methods to compute these anomalies were discussed in [40, 3, 45, 47, 46].
Note that the complete invertible symmetry of a UMTC is a 3-groupoid, $\underline{\underline{\text{Aut}(\mathcal{C})}}$. [^3] This is referred to as a categorical 2-group in [40]. SETs with symmetry group $G$ are completely distinguished by a 3-homomorphism $\underline{\underline{\rho}}:\underline{\underline{G}}\rightarrow\underline{\underline{\text{Aut}(\mathcal{C})}}$. See [40, 3, 48] for additional details.
@@ III.1 Review: gauged SPT defects
The above TQFT generally has finite invertible symmetries generated by operators called gauged SPT defects [24, 50]. These symmetry defects are defined by the SPT topological response with support on the submanifolds where the defects are located. Namely, the topological defect in an $n$-dimensional submanifold $M_{n}$ embedded in $(d+1)$D spacetime is defined as
@@
The gauged SPT defects also act on the junctions of the magnetic defects. For instance, in (2+1)D gauge theories suppose we have an operator corresponding to a network of magnetic fluxes. Then the gauged SPT defect acts non-trivially on the junctions of two magnetic fluxes $[g],[h]\in Cl(G)$ into $[gh]$, where $Cl(G)$ denotes the set of conjugacy classes of $G$. The 0-form symmetry generated by the (1+1)D gauged SPT defect acts on this junction by a phase factor $\exp(i\omega(g,h))$, since the SPT response evaluates nontrivially at this junction of $G$ defects. This implies that the symmetry operator acts on the junction of magnetic fluxes by a phase $\mathcal{U}_{\omega}$ given by
@@
Figure 4: The action of a gauged SPT defect on magnetic defects. (a): The 0-form symmetry generated by the gauged SPT defect acts on the magnetic defect by attaching the $(d-1)$D gauged SPT defect to the magnetic defect. This can be understood as the twisted compactification of SPT with respect to the circle with holonomy $g\in G$. (b): The 0-form gauged SPT defect in $(d+1)$D spacetime acts on a point junction of the $(d-1)$ magnetic defects by a phase factor given by the SPT cocycle. The figure shows the case of $d=2$, where the (1+1)D SPT defect acts on the junction of two magnetic defects by a 2-cocycle $\omega(g,h)$.
@@ III.2 Soft autoequivalence of finite gauge theory
When the braided fusion category is given by a (2+1)D finite gauge theory $\mathcal{C}=D(G)$, where $D(G)$ denotes the quantum double of $G$, a subgroup of the soft braided auto-equivalence $\mathrm{Aut}_{\mathrm{sf}}(\mathcal{C})$ has the structure of
$\mathrm{Out}_{\mathrm{sf}}(G)\ltimes H^{2}_{\mathrm{sf}}(BG,U(1))\subset\mathrm{Aut}_{\mathrm{sf}}(\mathcal{C})$ (31)
@@
where we label the anyon by the conjugacy class $[g_{j}]$ with $g_{j}\in{G_{\mathrm{sf}}}$ and the irreducible representation of the centralizer group $\pi_{j}\in\mathrm{Rep}(Z(g_{j}))$. Note that the above phase $\mathcal{U}_{pq}$ does not depend on the choice of representatives of the conjugacy class. We will later see that this symmetry does not permute anyons. As required by (12), the symmetry action leaves the $F$ symbol invariant, which is seen from the cocycle condition of group cohomology satisfied by $\mathcal{U}_{pq}$.
@@
In particular, when $g_{1}=p,g_{2}=q$ the $R$ symbol is transformed into $-R$, so it does not satisfy (13). Meanwhile, one can find another representative $\omega_{\text{sf}}\in Z^{2}(B{G_{\mathrm{sf}}},U(1))$ in the same cohomology class as $\omega_{pq}$, i.e., $[\omega_{\text{sf}}]=[\omega_{pq}]\in H^{2}(B{G_{\mathrm{sf}}},U(1))$ and symmetric as a function of $g,h$; $\omega_{\text{sf}}(g,h)=\omega_{\text{sf}}(h,g)$ for any $g,h\in{G_{\mathrm{sf}}}$. We will explicitly construct such a cocycle representative in Appendix A. With this new choice of 2-cocycle, the action on fusion vertices are given by
@@
and this phase action leaves both $F$, $R$ symbols invariant, being consistent with (12), (13). From now, we will use $U_{\text{sf}}$ as the data of the braided tensor auto-equivalence. This symmetry is regarded as the vertex basis gauge transformation with the matrices given by the phases $\Gamma=U_{\text{sf}}$. In other words, this ${\mathbb{Z}}_{2}$ symmetry is a soft symmetry. Below, let us summarize the notable properties of this ${\mathbb{Z}}_{2}$ symmetry:
@@
    which can be explicitly computed by using $\omega_{\text{sf}}$ presented in Appendix A. Since the above eigenvalue is invariant under the natural isomorphisms, this nontrivial invariant $(-1)$ implies that this soft ${\mathbb{Z}}_{2}$ symmetry corresponds to a nontrivial braided tensor auto-equivalence which cannot be given by natural isomorphisms. In Sec. IV, this ${\mathbb{Z}}_{2}$ symmetry will be realized in a non-Pauli stabilizer model of qubits. Therefore, this ${\mathbb{Z}}_{2}$ symmetry gives a non-trivial logical gate of the stabilizer code that does not permute anyons.
@@
While we discussed the soft symmetry in the untwisted ${G_{\mathrm{sf}}}$ gauge theory, this soft ${\mathbb{Z}}_{2}$ symmetry is present in ${G_{\mathrm{sf}}}$ gauge theory with a Dijkgraaf-Witten twist $\omega_{3}$ [49] as well, as long as the 2-cocycle $\omega_{\text{sf}}$ is not expressed as a slant product of $\omega_{\text{sf}}=i_{g}\omega_{3}$ with some $g\in{G_{\mathrm{sf}}}$. When $\omega_{\text{sf}}$ is the slant product, this SPT defect is bounded by the topological line operator of the magnetic flux. Concretely, suppose $\tilde{a}$ is the dual gauge field for $g\in{G_{\mathrm{sf}}}$, meaning $\exp(i\int\tilde{a})$ gives the line operator for the magnetic defect. We then have $d\tilde{a}=i_{g}\omega_{3}\mod 2\pi\mathbb{Z}$ [^4] More precisely, we have $d\tilde{a}=A^{*}(i_{g}\omega_{3})$ where $A:M_{3}\to B{G_{\mathrm{sf}}}$ is the background ${G_{\mathrm{sf}}}$ gauge field on the spacetime $M_{3}$. This means that the surface operator $\exp(i\int\omega_{\text{sf}})$ is bounded by the $g$ magnetic defect when $i_{g}\omega_{3}=\omega_{\text{sf}}\mod 2\pi\mathbb{Z}$. Since $\omega_{\text{sf}}$ is a coboundary, $\exp(i\int\omega_{\text{sf}})$ does not generate a faithful 0-form symmetry [50].
Also, we note that (2+1)D finite gauge theories can have a “softer” 0-form symmetry, which acts trivially on the Hilbert space of a genus $<g$ surface, but acts non-trivially on the genus $g$ surface. Such a global symmetry can be generated by the gauged SPT defects, where the SPT can only be detected by its partition function on higher genus surfaces. Such examples are explicitly constructed in Appendix B. It would be interesting to see if such softer symmetry can also arise from the outer automorphisms of gauge groups.

### v4

@@
(2+1)D topological orders possess emergent symmetries given by a group $\text{Aut}_{br}(\mathcal{C})$, which consists of the braided tensor autoequivalences of the modular tensor category $\mathcal{C}$ that describes the anyons. In this paper we discuss cases where $\text{Aut}_{br}(\mathcal{C})$ has elements that neither permute anyons nor are associated to any symmetry fractionalization but are still non-trivial, which we refer to as soft symmetries. We point out that one can construct topological defects corresponding to such exotic symmetry actions by decorating with a certain class of gauged SPT states that cannot be distinguished by their torus partition function. This gives a physical interpretation to work by Davydov on soft braided tensor autoequivalences. This has a number of important implications for the classification of gapped boundaries, non-invertible spontaneous symmetry breaking, and the general classification of symmetry-enriched topological phases of matter. We also demonstrate analogous phenomena in higher dimensions, such as (3+1)D gauge theory with gauge group given by the quaternion group $Q_{8}$.
@@ II.1 Review: Unitary modular tensor categories
where $\mu=1,\ldots,N_{ab}^{c}$, $d_{a}$ is the quantum dimension of $a$, and the factors $\left(\frac{d_{c}}{d_{a}d_{b}}\right)^{1/4}$ are a normalization convention for the diagrams. The set of anyons contains the identity or vacuum charge 0, and each anyon $a$ has a dual anti-particle $\bar{a}$ such that $a\times\bar{a}=0+\cdots$.
@@
[…]
  • $\varphi(a)=a$ for all anyons $a$.
  • The $U$ symbol $\{U_{\varphi}(a,b;c)\}$ is not a natural isomorphism.
@@ II.3 Symmetry action on genus $g$ surfaces
The $U$ symbols characterize the action of $\varphi]\in\text{Aut}_{br}(\mathcal{C})$ on the space of states on a closed genus $g$ surface. We can pick an orthonormal basis of states on the genus $g$ surface as
@@
It is straightforward to verify that in general the symmetry action on every basis state is unchanged by a natural isomorphism. Therefore, the symmetry action only depends on the equivalence class $\varphi]$. This implies that if a braided autoquivalence $\varphi$ acts as the identity on the space of states on a torus while it acts non-trivially on surfaces of genus $g\geq 2$, then it must be a non-trivial soft symmetry.
@@
The group of braided tensor autoequivalences defines an element of a 3rd cohomology group, $\mathcal{O}_{3}]\in\mathcal{H}^{3}(\text{Aut}_{br}(\mathcal{C}),\mathcal{A})$, where $\mathcal{A}$ is a finite Abelian group defined by fusion of the Abelian anyons in $\mathcal{C}$ [40, 3]. Formulas to compute $\mathcal{O}_{3}]$ from the $U$-symbols were presented in [3]. Graphically, we can understand $\mathcal{O}_{3}]$ as a failure of associativity for fusing the codimension-1 sheets. Consider three symmetries $\varphi_{1}]$, $\varphi_{2}]$, $\varphi_{3}]$ $\in\text{Aut}_{br}(\mathcal{C})$ and the junctions formed by fusing the sheets as $\varphi_{1}\circ(\varphi_{2}\circ\varphi_{3})$ compared with $(\varphi_{1}\circ\varphi_{2})\circ\varphi_{3}$. As we deform from the first order to the second, an Abelian anyon $O_{3}([\varphi_{1}],[\varphi_{2}],[\varphi_{3}])\in\mathcal{A}$ is emitted from the resulting transition point in space-time. See Fig. 2 for an illustration. We see how $\mathcal{O}_{3}]$ can be viewed as an obstruction to associativity of the symmetry operations [3].
The data $(\text{Aut}_{br}(\mathcal{C}),\mathcal{A},[\mathcal{O}_{3}])$ defines a 2-group $\underline{\text{Aut}_{br}(\mathcal{C})}$ that acts on the UMTC $\mathcal{C}$ [3, 41]. An example of a theory with non-trivial $\mathcal{O}_{3}]$ was discussed in [3, 42], with generalizations where the symmetry is anti-unitary [43].
@@ II.5 Symmetry fractionalization
where $\varphi_{jk}$ denotes $\varphi_{i}\circ\varphi_{j}$, and $M_{xa}$ denotes the $U(1)$ phase resulting from a double braid of $x$ with an Abelian anyon $a\in\mathcal{A}$. If $\mathcal{O}_{3}]$ is trivial in group cohomology $\mathcal{H}^{3}(\text{Aut}_{br}(\mathcal{C}),\mathcal{A})$ twisted by the action of $\text{Aut}_{br}(\mathcal{C})$ on $\mathcal{A}$, then there are inequivalent choices of the $\eta$ symbols, which form a torsor over the twisted cohomology $\mathcal{H}^{2}(\text{Aut}_{br}(\mathcal{C}),\mathcal{A})$.
The $\eta$-symbols are referred to as the symmetry fractionalization data of the theory, because they specify how the anyons carry fractional quantum numbers under the symmetry group elements. For example, they can be used to characterize the fractional electric charge of anyons in the FQH effect or fractional spin of spinons in quantum spin liquids [44]. Assuming the $\mathcal{O}_{3}]$ is trivial in $\mathcal{H}^{3}(\text{Aut}_{br}(\mathcal{C}),\mathcal{A})$, the $\eta$ symbols further determine an ’t Hooft anomaly $\mathcal{O}_{4}]\in\mathcal{H}^{4}(\text{Aut}_{br}(\mathcal{C}),U(1)) [$40, 3, 45, 46, 47].
@@ II.6 Global symmetry $G$ of quantum many-body system
Consider a microscopic quantum many-body system with a global 0-form symmetry group $G$. If the system forms a gapped phase with anyons described by a UMTC $\mathcal{C}$, then the possible symmetry-enriched topological phases will be partially characterized by a group homomorphism from the microscopic (UV) symmetry to the emergent (IR) symmetry,
$[\rho]:G\rightarrow\text{Aut}_{br}(\mathcal{C}).$ (21)
@@
The group homomorphism $\rho]$ can be pulled back to determine a cohomology class $\rho]^{*}[\mathcal{O}_{3}]\in\mathcal{H}^{3}(G,\mathcal{A})$, which is an obstruction for the $\eta$ symbols of $G$ to satisfy associativity. Let us assume that $\rho]^{*}[\mathcal{O}_{3}]$ is trivial.
@@
[…]
$\underline{\rho}:\underline{G}\rightarrow\underline{\text{Aut}_{br}(\mathcal{C})}.$ (24)
If $\rho]^{*}[\mathcal{O}_{3}]$ is non-trivial, then the theory is incompatible with an ordinary global $G$ symmetry. Instead we must consider a system with a non-trivial 2-group symmetry $\underline{G}=(G,\mathcal{A},[\rho],[\rho]^{*}[\mathcal{O}_{3}])$. In this sense $\rho]^{*}[\mathcal{O}_{3}]$ is an obstruction to having the symmetry be an ordinary group $G$.
As mentioned above, the symmetry fractionalization data determines an element $\mathcal{O}_{4}]\in\mathcal{H}^{4}(\text{Aut}_{br}(\mathcal{C}),U(1))$, which in the field theory literature is referred to as the ’t Hooft anomaly. This induces an ’t Hooft anomaly for $G$ via the pullback $\mathcal{O}_{4,G}]=[\rho]^{*}[\mathcal{O}_{4}]\times\mathcal{O}_{4,\text{rel}}[{\bf t}]$, where $\mathcal{O}_{4,\text{rel}}[{\bf t}]$ is an additional contribution, referred to as a relative anomaly [46], determined by the choice $\bf t]$, along with the $\eta$ and $U$ symbols for $\text{Aut}_{br}(\mathcal{C})$. General methods to compute these anomalies were discussed in [40, 3, 45, 47, 46].
Note that the complete invertible symmetry of a UMTC is a 3-groupoid, $\underline{\underline{\text{Aut}_{br}(\mathcal{C})}}$. [^3] This is referred to as a categorical 2-group in [40]. SETs with symmetry group $G$ are completely distinguished by a 3-homomorphism $\underline{\underline{\rho}}:\underline{\underline{G}}\rightarrow\underline{\underline{\text{Aut}_{br}(\mathcal{C})}}$. See [40, 3, 48] for additional details.
@@ III.1 Review: gauged SPT defects
The above TQFT generally has finite invertible symmetries generated by operators called gauged SPT defects [24, 50]. These symmetry defects are defined by the SPT topological response with support on the submanifolds where the defects are located. Namely, the topological defect in an $n$-dimensional submanifold $M_{n}$ embedded in (d+1)(d+1)D spacetime is defined as
@@
The gauged SPT defects also act on the junctions of the magnetic defects. For instance, in (2+1)D gauge theories suppose we have an operator corresponding to a network of magnetic fluxes. Then the gauged SPT defect acts non-trivially on the junctions of two magnetic fluxes $g],[h]\in Cl(G)$ into $[gh]$, where $Cl(G)$ denotes the set of conjugacy classes of $G$. The 0-form symmetry generated by the (1+1)D gauged SPT defect acts on this junction by a phase factor $\exp(i\omega(g,h))$, since the SPT response evaluates nontrivially at this junction of $G$ defects. This implies that the symmetry operator acts on the junction of magnetic fluxes by a phase $\mathcal{U}_{\omega}$ given by
@@
Figure 4: The action of a gauged SPT defect on magnetic defects. (a): The 0-form symmetry generated by the gauged SPT defect acts on the magnetic defect by attaching the (d1)(d-1)D gauged SPT defect to the magnetic defect. This can be understood as the twisted compactification of SPT with respect to the circle with holonomy $g\in G$. (b): The 0-form gauged SPT defect in (d+1)(d+1)D spacetime acts on a point junction of the (d1)(d-1) magnetic defects by a phase factor given by the SPT cocycle. The figure shows the case of $d=2$, where the (1+1)D SPT defect acts on the junction of two magnetic defects by a 2-cocycle $\omega(g,h)$.
@@ III.2 Soft autoequivalence of finite gauge theory
Suppose that the braided fusion category is given by a (2+1)D finite gauge theory $\mathcal{C}=Z(\text{Vec}(G))$, where $Z(\text{Vec}(G))$ denotes the Drinfeld center of $\text{Vec}(G)$. Its simple anyons are labeled by a pair
$([g],\pi)~,$ (30)
where $g]\in Cl(G)$ is a conjugacy class of $G$ that specifies the magnetic flux, and $\pi\in\text{Rep}(Z(G))$ is the irreducible representation that corresponds to the electric charge.
Then, a subgroup of the soft braided auto-equivalence $\mathrm{Aut}_{\mathrm{sf}}(\mathcal{C})$ has the structure of
$\mathrm{Out}_{\mathrm{sf}}(G)\ltimes H^{2}_{\mathrm{sf}}(BG,U(1))\subset\mathrm{Aut}_{\mathrm{sf}}(\mathcal{C})~.$ (31)
@@
where we label the anyon by the conjugacy class $g_{j}]$ with $g_{j}\in{G_{\mathrm{sf}}}$ and the irreducible representation of the centralizer group $\pi_{j}\in\mathrm{Rep}(Z(g_{j}))$. Note that the above phase $\mathcal{U}_{pq}$ does not depend on the choice of representatives of the conjugacy class. We will later see that this symmetry does not permute anyons. As required by (11), the symmetry action leaves the $F$ symbol invariant, which is seen from the cocycle condition of group cohomology satisfied by $\mathcal{U}_{pq}$.
@@
In particular, when $g_{1}=p,g_{2}=q$ the $R$ symbol is transformed into R-R, so it does not satisfy (12). Meanwhile, one can find another representative $\omega_{\text{sf}}\in Z^{2}(B{G_{\mathrm{sf}}},U(1))$ in the same cohomology class as $\omega_{pq}$, i.e., $\omega_{\text{sf}}]=[\omega_{pq}]\in H^{2}(B{G_{\mathrm{sf}}},U(1))$ and symmetric as a function of $g,h$; $\omega_{\text{sf}}(g,h)=\omega_{\text{sf}}(h,g)$ for any $g,h\in{G_{\mathrm{sf}}}$. We will explicitly construct such a cocycle representative in Appendix A. With this new choice of 2-cocycle, the action on fusion vertices are given by
@@
and this phase action leaves both $F$, $R$ symbols invariant, being consistent with (11), (12). From now, we will use $U_{\text{sf}}$ as the data of the braided tensor auto-equivalence. This symmetry is regarded as the vertex basis gauge transformation with the matrices given by the phases $\Gamma=U_{\text{sf}}$. The symmetry does not permute anyons, while as demonstrated below, this vertex gauge transformation is not a natural isomorphism. In other words, this ${\mathbb{Z}}_{2}$ symmetry is a soft symmetry. Below, let us summarize the notable properties of this ${\mathbb{Z}}_{2}$ symmetry:
@@
    which can be explicitly computed by using $\omega_{\text{sf}}$ presented in Appendix A. Since the above eigenvalue is invariant under the natural isomorphisms, this nontrivial invariant (1)(-1) implies that this soft ${\mathbb{Z}}_{2}$ symmetry corresponds to a nontrivial braided tensor auto-equivalence which cannot be given by natural isomorphisms. In Sec. IV, this ${\mathbb{Z}}_{2}$ symmetry will be realized in a non-Pauli stabilizer model of qubits. Therefore, this ${\mathbb{Z}}_{2}$ symmetry gives a non-trivial logical gate of the stabilizer code that does not permute anyons.
@@
[…]
Also, we note that (2+1)D finite gauge theories can have a “softer” 0-form symmetry, which acts trivially on the Hilbert space of a genus <g<g surface, but acts non-trivially on the genus $g$ surface. Such a global symmetry can be generated by the gauged SPT defects, where the SPT can only be detected by its partition function on higher genus surfaces. Such examples are explicitly constructed in Appendix B. It would be interesting to see if such softer symmetry can also arise from the outer automorphisms of gauge groups.